OpenStudy (anonymous):

r=4q-5p/9 solve for p

OpenStudy (anonymous):

-5p/9=r-4q -5p=9r-16q p=-9/5 r +16/5 q

OpenStudy (anonymous):

that is very confusing! lol i am not doing very good on this final. i still have at least ten more problems dealing with slope intercept equations.. thanks for your help!